# Geometric Optimal Control of Mechanical Systems with Gravitational and Resistive Force

### v1

@@ Abstract
Optimal control plays a crucial role in numerous mechanical and robotic applications. Broadly, optimal control methods are divided into direct methods (which optimize trajectories directly via discretization) and indirect methods (which transform optimality conditions into equations that guarantee optimal trajectories). While direct methods could mask geometric insights into system dynamics due to discretization, indirect methods offer a deeper understanding of the system’s geometry. In this paper, we propose a geometric framework for understanding optimal control in mechanical systems, focusing on the combined effects of inertia, drag, and gravitational forces. By modeling mechanical systems as configuration manifolds equipped with kinetic and drag metrics, alongside a potential field, we explore how these factors influence trajectory optimization. We derive optimal control equations incorporating these effects and apply them to two-link and UR5 robotic manipulators, demonstrating how manifold curvature and resistive forces shape optimal trajectories. This work offers a comprehensive geometric approach to optimal control, with broad applications to robotic systems.
@@ I INTRODUCTION
The optimal control has a long and rich history, from the Brachistochrone curve to various applications in robotics, such as manipulators, legged locomotion systems [1, 2, 3], and undulatory locomotion systems [4, 5, 6]. Optimal control methods can generally be categorized into two main approaches: direct and indirect methods.
Direct methods involve discretizing trajectories and using the waypoint locations (or other parameterization) as decision variables to solve the optimal control problem directly. Although this approach is relatively straightforward to implement, it can obscure the geometric properties of the problem. In contrast, indirect methods employ the Pontryagin Maximum Principle to transform the optimal control problem into a differential equation. Then, a problem structure of the optimal control can be seen as a boundary value problem, which can be solved via a shooting method. [^1] This method can be implemented by these steps: 1) fix some initial conditions, 2) iterate over high-order initial conditions while shooting guaranteed-optimal trajectories to hit low-order final conditions. These differential equations often offer deeper geometric insights into the optimal trajectories. For this paper, we focus on geometric optimal control, which primarily focuses on indirect methods.
@@
Figure 1: (a) Optimal trajectories for a UR5 moving its first three joints. The black-$\times$-marked curve denotes the optimal trajectory considering only the inertia effect. The distortion based on the curvature of the inertia manifolds adds acceleration to the trajectory in the direction to reduce the inertia effect (e.g., the configuration folding the arm). The red-$\circ$-marked curve includes both inertia and gravitational effects, influenced by the curvature of the potential field on the inertia manifold. The blue curve denotes the optimal trajectory considering both inertia and drag, where drag force is modeled as a joint viscous friction. In this case, a drag metric reduces to an Euclidean metric, shortening the optimal trajectory in configuration space. (b) The same optimal trajectories in task space as those in configuration space with the initial and final configuration of UR5 manipulators. Note that we only give the command to the first three joints starting from the base joint for this work.
In Lagrangian mechanics, the mechanical systems can be expressed in terms of a Riemannian manifold associated with a metric [7, 8]. The Riemannian metric in this context represents the kinetic energy as a function of the system velocity. Previous works [9, 10] have derived optimal control equations using the Pontryagin Maximum Principle under acceleration- or torque-based costs. These approaches focus on minimizing the magnitude of accelerations on inertially defined Riemannian manifolds. Additionally, the influence of potential fields on optimal control equations has been considered through the Riemannian Hessian of the potential field [11, 12].
Similarly, drag-dominated mechanical systems are associated with a Riemannian manifold equipped with a drag metric tensor [13, 5]. This drag metric tensor represents the drag coefficients at each configuration and describes the power dissipated by configuration velocities. In such systems, geodesics on drag manifolds represent the solutions to optimal control problems, as these geodesics find the shortest path and constant-speed profile on the drag manifolds. However, little work has been done on the geometric understanding of optimal trajectories considering the combined effects of kinetic, drag, and gravitational forces beyond these two extreme cases.
In this paper, we aim to provide a geometric understanding of optimal trajectories considering the influence of various factors and derive an optimal control equation with a general force expressed as a function of configuration and velocity. As illustrated in Fig. 1, we identify three key elements affecting optimal trajectories:
@@
    Inertial Effect: Optimal trajectories are influenced by the curvature of the inertia manifolds. Generally, positive curvature tends to make optimal trajectories converge at specific points [10]. To minimize acceleration- or torque-norm costs, the trajectory on the positive curvature of manifolds generates the acceleration in the direction pointing inward.
@@
    Potential Force Effect: Similar to the inertial effect, optimal trajectories are affected by the curvature of potential manifolds, which are associated with the inertia manifolds [11, 12]. These manifolds are essentially the graphs of potential energy plotted on the inertia manifolds.
@@
    Resistive Force Effect: This effect tends to align optimal trajectories with the drag manifolds. To minimize energy dissipation due to resistive forces, such as joint friction or end-effector drag, it is preferable to follow the shortest path on drag manifolds [5].
To demonstrate our approach and provide a geometric understanding, we apply the optimal control equation to two-link robotic manipulators with rigid bodies and the UR5 manipulator, considering each effect seperately in section IV and combining effect in section V. The two-link robotic manipulator, defined by two joint variables, is particularly useful for illustrating the associated manifolds, such as potential fields and drag manifolds. The UR5 is a widely used 6-DOF robotic manipulator in research and industrial applications [14]. In this robot, the first three joints control the position of the end-effector, while the last three joints control its orientation. To highlight the effects of inertia and gravity, we focus on the optimal trajectories within the reduced configuration space defined by the first three joints.
@@ II-A Riemmanian geometry
Figure 2: A one-parameter family of optimal trajectories in inertia manifolds of different curvatures. $s$ is a variation parameter. $t$ is a time parameter of each curve. The vector field $u$ is a control input. Black curves denote one of the optimal trajectories in the family. (a) Optimal trajectories families in flat manifolds. Because of the zero curvature, the optimal trajectory is a fourth-order polynomial spline equation [10]. (b) Optimal trajectories families in manifolds of positive curvature (e.g., spherical type of manifolds). Because of positive curvature, a direction of control inputs $u$ is gradually turning inward. (c) Optimal trajectories families in manifolds of negative curvature (e.g., saddle type of manifolds). Because of positive curvature, a direction of control inputs $u$ is gradually turning outward.
A Riemannian metric $M$ for the space contains the distance information, and a manifold $Q$ equipped with a metric tensor $M$ is called a Riemannian manifold $(Q,M)$. Let $U\in\mathbb{R}^{d}$ be an open subset with coordinates ${q}=({q}^{1},\cdots,{q}^{d}).$ $U$ is a coordinate domain in a configuration manifold $Q$, and it allows to express explicit coordinate formulas for the tensors. ($\partial_{{q}^{i}}\equiv\frac{\partial}{\partial{q}^{i}}$) denotes the basis of $TU$ given by (constant) vector fields in $U$ and ($\partial{q}^{i}$) the dual basis of $T^{*}U$.
A (tangent) vector on a manifold describes the rate at which a point moves through a Riemannian manifold $Q$. A covector describes the rate where a quantity changes across a space at a given point. A covector has a natural product with a tangent vector because it acts as a linear map from tangent vectors to scalars. It is called a covector-vector product,
@@
$\left\langle a,\bullet\right\rangle_{M}$ $=M_{ij}a^{i}\partial{q}^{j}=F,$ (2b)
where $\left\langle\cdot,\cdot\right\rangle_{M}$ is the inner product with respect to $M$, $F\in T_{{q}}^{*}Q$ is a covector, $a\in T_{{q}}Q$ is a tangent vector, and $M_{ij}$, the bullet notations $\bullet$ represent an element of $T_{{q}}Q$ or $T_{{q}}^{*}Q$ so that the corresponding linear function is a dual element in $T_{{q}}^{*}Q$ or $T_{{q}}Q$, and $M^{ij}$ are the components of $M$ and its inverse (i.e., dual), respectively. The operations in 2a and 2b raise and lower vector indices, transforming between vectors and covectors. [^3] These operations are known as a musical isomorphism operator. For raising an index, a musical sharp symbol is used ($a=F^{\sharp}$). For lowering an index, a musical flat symbol is used ($F=a^{\flat}$). From the viewpoint of mechanics, raising the index transforms a force covector into an acceleration vector through the mass matrix $M({q})$.
Unlike a flat space, the local coordinate bases of a curved space can vary at each point. A covariant derivative is a coordinate-invariant way to differentiate tensor (e.g., vector or covector) fields along curves by including correction terms for changes in the local coordinate bases. A covariant derivative of a vector $\nabla_{\dot{{q}}}\dot{{q}}$ and a covector $\nabla_{\dot{{q}}}^{*}F$ along the curve ${q}(t)$ can be expressed in natural coordinates induced by coordinates $({q}^{i})$ in $Q$:
@@
where $\dot{{q}}(t)$ is a time-parameterized curve, $\nabla$ and $\nabla^{*}$ are the Levi-Civita Connection corresponding to a metric $M$ and its dual connection, $\Gamma^{i}_{jk}$ is Christoffel symbols [15]. A covariant acceleration $\nabla_{\dot{{q}}}\dot{{q}}$ is a time-covariant derivative of a velocity $\dot{{q}}$ along a curve ${q}(t)$. If a covariant acceleration is zero, ${q}(t)$ is called a geodesic.
A Riemannian curvature tensor $R\in\Omega^{2}(Q,End(TQ))$ describes a curvature of the curvature of Riemannian manifolds and is defined by a commutator of a covariant derivative,
$R(X,Y)Z=\nabla_{X}\nabla_{Y}Z-\nabla_{Y}\nabla_{X}Z-\nabla_{\left[X,Y\right]}Z,$ (5)
[…]
$\left\langle\partial{q}^{l},R(\partial_{{q}^{i}},\partial_{{q}^{j}})\partial_{{q}^{k}}\right\rangle=R^{l}_{ijk}({q})\\ =\partial_{{q}^{i}}\Gamma^{l}_{jk}-\partial_{{q}^{j}}\Gamma^{l}_{jk}+\Gamma^{n}_{jk}\Gamma^{l}_{in}-\Gamma^{n}_{ik}\Gamma^{l}_{jn},$ (6)
@@ III-A Optimal Control System
The optimal control system is defined by the data ($Q,M,N,a_{\text{ext}}$), where $Q$ is configuration manifold, $M$ is a “mass” metric, $N$ is an induced metric encoding control cost [10], and $a_{\text{ext}}=\left\langle F_{\text{ext}},\bullet\right\rangle_{M^{*}}$ is an equivalent acceleration field generated by a external force $F_{\text{ext}}$ dependent on configuration and velocity, including gravitational [^4] Although a Lagrangian includes potential energy, we interpret a gravitational force as an external force to “inertia manifold” in this paper. and drag forces.
The system is given by:
@@
Initial Condition: ${q}(0)={q}_{0},\ \ \dot{{q}}(0)=v_{0},$ (7d)
Final Condition: ${q}(T)={q}_{T},\ \dot{{q}}(T)=v_{T},$ (7e)
where $u$ is a control input, and $T$ is a trajectory time. Raising an index in the control equation gives the Euler-Lagrange equation,
$M({q})\ddot{{q}}+C({q},\dot{{q}})=M({q})u+F_{\text{ext}}({q},\dot{{q}}),$ (8)
where $C({q},\dot{{q}})$ represents centrifugal and Coriolis forces, interpretable as Christoffel symbols. We use an acceleration-based input norm with respect to $N$ for the cost functional. Based on the induced metric, a cost functional has a different meaning. For example, the actuator-torque norm cost can be defined by
@@
with $\tilde{M}$ being the torque-based cometric tensor, which reduces to the Euclidean metric when joint and actuator configurations are identical [10].
@@ III-B Optimal Spline Equation
Figure 3: (a) A two-link manipulator and optimal trajectories under situations with and without gravitation force in task space. The mass and the length of each link are the same ($m,\ell$). ${q}_{i}$ is a joint angle of $i$-th link and a configuration. The direction of gravity force is toward the bottom. (b) The same optimal trajectories in configuration space. The potential field is visualized via a filled contour. The inertia metric is visualized by their Tissot’s indicatrix (ellipse and cross lines) which describes how much the space shrinks or expands at each point in a geometric viewpoint or how much velocity is needed at each point to produce unit kinetic energy in a mechanics viewpoint. (c) Optimal trajectories on a potential field with a metric-stretched space. The metric-stretched space approximates the geometry of the manifold such that basis vectors in each direction at each point correspond to the same kinetic energy as much as possible. This algorithm is presented in [16]. The blue- and black-colored curves respectively are the optimal trajectories generated by the optimal control of the two-link manipulator with and without gravity. In this space, the Tissot indicatrix ellipses have a circular-like shape. A surface at the top visualizes the potential field. The normal distance between the stretched surface and each point on the potential field surface represents potential energy at each configuration. We use the same colormap of a filled contour as that of (b).
@@
$H=T^{*}Q\times_{Q}TQ\times_{Q}T^{*}Q,\ \Omega=\omega_{\nabla},\\ \text{ with variables }p,\alpha\in T_{{q}}^{*}Q,v\in T_{{q}}^{*}Q,$ (10)
@@
$H_{u}=-\lVert u\rVert_{N}^{2}+\left\langle p,v\right\rangle+\left\langle\alpha,u+a_{\text{ext}}\right\rangle,$ (11)
@@
$H=\frac{1}{4}\lVert\alpha\rVert_{N^{*}}^{2}+\left\langle\alpha,a_{\text{ext}}\right\rangle+\left\langle p,v\right\rangle,\ u_{max}=\dfrac{1}{2}N^{*}\alpha.$ (12)
Using the formulas for $\omega_{\nabla}$ and the corresponding Hamiltonian vector fields in [9, Prop. 1], we arrive at the following system of equations written in natural coordinates induced by coordinates $({q}^{i})$ in $Q$,
@@
$\dot{{q}}^{i}$ $=v^{i},$
$\left(\nabla_{v}v\right)^{i}$ $=\frac{1}{2}N^{ij}\alpha_{j}+a_{\text{ext}}^{i},$
$\left(\nabla_{v}^{*}\alpha\right)_{i}$ $=-p_{i}-\alpha_{j}(\partial_{v^{i}}a_{\text{ext}}^{j}),$
$\left(\nabla_{v}^{*}p\right)_{i}$ $=-\frac{1}{4}\tau_{i}^{jk}\alpha_{j}\alpha_{k}+R_{ijk}^{l}v_{j}v_{k}\alpha_{l}$
$-\big{[}\underbrace{\alpha_{k}\left(\nabla_{\partial_{{q}^{i}}}a_{\text{ext}}\right)^{k}-\Gamma_{ij}^{k}v^{j}(\partial_{v^{k}}a_{\text{ext}}^{l})\alpha_{l}}_{\text{External force terms}}\big{]},$ (13)
where, as in [10, Appendix A], $R$ and $\nabla$ are associated with a metric tensor $M$, and the tensor $\tau_{i}^{jk}\equiv\nabla_{\bullet}^{*}N^{*}$ is given by a covariant derivative of a dual induced tensor,
@@
We skip the coordinate-invariant version of the equations for the purpose of this paper. The coordinate-invariant approach will be discussed elsewhere.

### v2

# Riemannian Variational Calculus: Optimal Trajectories Under Inertia, Gravity, and Drag Effects

@@ Abstract
Robotic motion optimization often focuses on task-specific solutions, overlooking fundamental motion principles. Building on Riemannian geometry and the calculus of variations (often appearing as indirect methods of optimal control), we derive an optimal control equation that expresses general forces as functions of configuration and velocity, revealing how inertia, gravity, and drag shape optimal trajectories. Our analysis identifies three key effects: (i) curvature effects of inertia manifold, (ii) curvature effects of potential field, and (iii) shortening effects from resistive force. We validate our approach on a two-link manipulator and a UR5, demonstrating a unified geometric framework for understanding optimal trajectories beyond geodesic-based planning.
@@ I INTRODUCTION
Recent advances in computational power and numerical control algorithms have enabled robots (or machines) to execute complex, precise motions across a wide range of applications. In particular, numerical trajectory optimization methods (appearing as direct methods of optimal control [1]) are widely used to generate task-specific solutions, and learning-based control methods [2] allow robots to generate efficient motions without explicit models by learning from input-output data. However, these approaches often focus on optimizing individual motion strategies without explicitly characterizing the fundamental motion principles—such as how optimal motion strategies adapt to external forces or kinetic energy changes—thereby hindering the development of more generalizable and interpretable motion strategies.
On the other hand, the calculus of variations (appearing as indirect methods of optimal control [3, 4, 5]) provides a systematic framework for understanding how small variations in trajectories affect optimality. In Riemannian geometry, it extends to describe a system’s optimal dynamics, by setting a Riemannian metric encoding the system’s kinetic energy or power dissipated by environments [6, 7, 8, 9]. In this setting, geodesics—the shortest or energy-optimal paths on the manifold—naturally emerge as optimal motion strategies, governed by the geodesic equation, a system of differential equations that characterizes these paths.
@@
Figure 1: (a) Optimal trajectories for a UR5 manipulator moving its first three joints. The black-$\times$-marked curve considers only inertia effects, where curvature-induced distortion directs acceleration to reduce inertia (e.g., folding the arm). The red-$\circ$-marked curve incorporates both inertia and gravity, influenced by the potential field curvature on the inertia manifold. The blue curve accounts for inertia and drag, where drag, modeled as joint viscous friction, results in an Euclidean drag metric, shortening the trajectory in configuration space. (b) The same optimal trajectories in task space, with the initial and final configurations of the UR5. Only the first three joints, starting from the base joint, are actuated in this work. The manipulator’s appearance transitions from white to gray, representing its initial and final configurations.
Although the geodesic equation serves as the optimal control equation for kinematic systems and drag-dominated systems [9], it does not account for second-order dynamics. For mechanical systems without friction, previous works [3, 4, 10, 11] have introduced the Riemannian spline equation to characterize natural (optimal) motion by minimizing acceleration. Additionally, the influence of potential fields on the Riemannian spline equation has been considered [3, 10]. Our recent studies [12, 13] introduce an actuation-based cometric and propose a biased Riemannian spline equation to minimize actuation effort. However, to the best of our knowledge, the geometric understanding of optimal trajectories considering the combined effects of kinetic, drag, and gravitational forces remains largely unexplored.
In this paper, building on previous works on Riemannian splines [11, 12, 13], we provide a geometric interpretation of optimal trajectories under the combined effects of general forces, in which a general force is expressed as a function of configuration (e.g., joint angles or shape) and velocity. As illustrated in Fig. 1, three key elements shape the structure of optimal trajectories:
@@
    Inertial Effect: Optimal trajectories are influenced by the curvature of the inertia manifold. Generally, positive curvature tends to make trajectories converge at specific points [12].
@@
    Potential Force Effect: Similar to inertia effects, optimal trajectories are affected by the curvature of potential manifolds, which are linked to inertia manifolds [3, 10]. These manifolds represent graphs of potential energy over inertia manifolds.
@@
    Resistive Force Effect: This effect aligns optimal trajectories with drag manifolds. To minimize energy dissipation from resistive forces, such as joint friction or end-effector drag, it is preferable to follow the shortest path on drag manifolds [9].
To demonstrate our approach and provide a geometric interpretation, we apply the optimal control equation to a two-link manipulator and a UR5 manipulator, analyzing individual effects in section IV and their combined influence in section V. The two-link manipulator, with two joint variables, effectively illustrates key manifolds such as potential fields and drag manifolds. The UR5, a widely used 6-DOF robotic manipulator in research and industry [14], has its first three joints governing end-effector position and the last three controlling orientation. To emphasize inertia and gravity effects, we focus on optimal trajectories in the reduced configuration space of the first three joints.
@@ II-A Riemmanian geometry
A locally differentiable manifold $Q$ equipped with a metric tensor $M$ is called a Riemannian manifold $(Q,M)$. Let $U\in\mathbb{R}^{d}$ be an open subset with coordinates ${q}=({q}^{1},\cdots,{q}^{d}).$ $M(q)$ is a Riemannian metric for the space contains the local distance information. $U$ is a coordinate domain in a configuration manifold $Q$, and it allows to express explicit coordinate formulas for the tensors. ($\partial_{{q}^{i}}\equiv\frac{\partial}{\partial{q}^{i}}$) denotes the basis of $TU$ given by (constant) vector fields in $U$ and ($\partial{q}^{i}$) the dual basis of $T^{*}U$.
A tensor is a mathematical object that generalizes scalars, vectors, and matrices to higher dimensions while maintaining certain transformation properties under coordinate changes. A tensor of type $(m,n)$ is a multilinear map that takes in $m$ covectors (covariant components) and $n$ vectors (contravariant components) and produces a scalar [15]. In coordinates, it is typically represented by a multi-indexed array of numbers. [^1] We use an Einstein summation convention. A lower index indicates covector elements and an upper index indicates vector elements. A (tangent) vector on a manifold, as a $(1,0)$-tensor, describes the rate at which a point moves through a Riemannian manifold $Q$. A covector, as a $(0,1)$-tensor, describes the rate where a quantity changes across a space at a given point. There is a natural product between a covector and a vector because a covector acts as a linear map from tangent vectors to scalars. It is called a covector-vector product,
@@
$\left\langle\bullet,a\right\rangle_{M}$ $=M_{ij}a^{i}\partial{q}^{j}=F,$ (2b)
where $\left\langle\cdot,\cdot\right\rangle_{M}$ is the inner product with respect to $M$, $F\in T_{{q}}^{*}Q$ is a covector, $a\in T_{{q}}Q$ is a tangent vector, the bullet notations $\bullet$ represent an element of $T_{{q}}Q$ or $T_{{q}}^{*}Q$ so that the corresponding linear function is a dual element in $T_{{q}}^{*}Q$ or $T_{{q}}Q$, $M_{ij}$ are the components of $M$, and $M^{ij}$ are the components of the metric inverse (i.e., dual of metric) $M^{*}$, respectively. The operations in 2a and 2b raise and lower vector indices, transforming between vectors and covectors. [^2] These operations are known as a musical isomorphism operator. For raising an index, a musical sharp symbol is used ($a=F^{\sharp}$). For lowering an index, a musical flat symbol is used ($F=a^{\flat}$). From the viewpoint of mechanics, raising the index transforms a force covector into an acceleration vector through the mass matrix $M({q})$ [16].
A covariant derivative is a way to differentiate vector fields (or tensor fields) on a curved space or manifold while respecting the manifold’s geometry. Unlike the standard derivative, which does not account for changes in the coordinate basis, the covariant derivative subtracts out the “expected changes” in the fields induced by “warping” of the coordinate chart with respect to the manifold, and so ensures that the result only includes “true changes” to the field across the manifold. Given a smooth manifold $Q$ with a Levi-Civita connection $\nabla$ and its dual connection $\nabla^{*}$, a covariant derivative of a vector ${q}$ and a covector $F$ along the curve ${q}(t)$ can be expressed in natural coordinates induced by coordinates $({q}^{i})$ in $Q$:
@@
where ${q}(t)$ is a time-parameterized curve, and $\Gamma^{i}_{jk}$ is Christoffel symbols [15]. A covariant acceleration $\nabla_{\dot{{q}}}\dot{{q}}$ is a covariant time-derivative of a velocity $\dot{{q}}$ along a curve ${q}(t)$. If a covariant acceleration is zero, ${q}(t)$ is called a geodesic.
A Riemannian curvature tensor $R\in\Omega^{2}(Q,End(TQ))$ (see [11]) describes a curvature of the curvature of Riemannian manifolds and its coefficient function $R^{l}_{ijk}({q})$ is defined by
$R^{l}_{ijk}({q})=\partial_{{q}^{i}}\Gamma^{l}_{jk}-\partial_{{q}^{j}}\Gamma^{l}_{jk}+\Gamma^{n}_{jk}\Gamma^{l}_{in}-\Gamma^{n}_{ik}\Gamma^{l}_{jn}.$ (5)
@@ III-A Optimal Control System
The optimal control system is defined by the data ($Q,M,N,F_{\text{ext}}$), where $Q$ is configuration manifold, $M$ is a “mass” metric, $N$ is an induced metric encoding control cost [12], and $F_{\text{ext}}$ is an external force dependent on configuration and velocity, including gravitational [^3] Although a Lagrangian includes potential energy, we interpret a gravitational force as an external force to “inertia manifold” in this paper. and drag forces.
The system is defined as follows:
@@
Initial Condition: ${q}(t_{0})={q}_{0},\ \ \dot{{q}}(t_{0})=v_{0},$ (6d)
Final Condition: ${q}(t_{f})={q}_{f},\ \dot{{q}}(t_{f})=v_{f},$ (6e)
where $u$ is a control input, and $t_{0}$ and $t_{f}$ denote the initial and final times of trajectory, respectively. The term $a_{\text{ext}}=\left\langle F_{\text{ext}},\bullet\right\rangle_{M^{*}}$ represents the acceleration field induced by external force $F_{\text{ext}}$. Note that the control equation (6b) corresponds to the system’s equation of motion where the mass matrix is given by the metric tensor $M$. The interpretation of the cost functional depends on the induced metric. If the induced metric coincides with the dual metric, the cost functional represents an acceleration-based input norm. Given the actuation-based cometric, the induced metric and the actuator-torque norm cost can be defined as
@@
with $\tilde{M}$ being the actuation-based cometric tensor, which simplifies to the Euclidean metric when joint and actuator configurations are identical (see [12, 13]).
@@ III-B Optimal Spline Equation
Figure 2: (a) A two-link manipulator and its optimal trajectories in task space, with and without gravity. Each link has the same mass ($m$) and length ($\ell$), and the joint angle of the $i$-th link is denoted as ${q}_{i}$. Gravity acts downward. (b) The same optimal trajectories in configuration space, where the potential field is shown as a filled contour. Tissot’s indicatrix (ellipses and cross lines) visualizes the inertia metric, indicating spatial contraction/expansion and the velocity needed for unit kinetic energy. (c) Optimal trajectories in a metric-stretched space, approximating the Riemannian manifold’s geometry by scaling basis vectors for uniform kinetic energy representation [9].
@@
$H=T^{*}Q\times_{Q}TQ\times_{Q}T^{*}Q,\ \Omega=\omega_{\nabla},\\ \text{ with variables }p,\alpha\in T_{{q}}^{*}Q,\dot{{q}}\in T_{{q}}^{*}Q,$ (8)
@@
$H_{u}=-N_{ij}{u}^{i}{u}^{j}+p_{i}\dot{{q}}^{i}+\alpha_{i}u^{i}+\alpha_{i}a_{\text{ext}}^{i},$ (9)
@@
$H=\frac{1}{4}N^{ij}\alpha_{i}\alpha_{j}+\alpha_{i}a_{\text{ext}}^{i}+p_{i}\dot{{q}}^{i},\ (u_{max})^{i}=\dfrac{1}{2}N^{ij}\alpha_{j}.$ (10)
Using the formulas for $\omega_{\nabla}$ and the corresponding Hamiltonian vector fields in [11, Prop. 1], we arrive at the following system of equations written in natural coordinates induced by coordinates $({q}^{i})$ in $Q$,
@@
$\left(\nabla_{\dot{{q}}}\dot{{q}}\right)^{i}$ $=\frac{1}{2}N^{ij}\alpha_{j}+a_{\text{ext}}^{i},$
$\left(\nabla_{\dot{{q}}}^{*}\alpha\right)_{i}$ $=-p_{i}-\alpha_{j}(\partial_{\dot{{q}}^{i}}a_{\text{ext}}^{j}),$
$\left(\nabla_{\dot{{q}}}^{*}p\right)_{i}$ $=-\frac{1}{4}\tau_{i}^{jk}\alpha_{j}\alpha_{k}+R_{ijk}^{l}\dot{{q}}^{j}\dot{{q}}^{k}\alpha_{l}$
$-\big{[}\underbrace{\alpha_{k}\left(\nabla_{\partial_{{q}^{i}}}a_{\text{ext}}\right)^{k}-\Gamma_{ij}^{k}\dot{{q}}^{j}(\partial_{\dot{{q}}^{k}}a_{\text{ext}}^{l})\alpha_{l}}_{\text{External force terms}}\big{]},$ (11)
where, as in [12, Appendix A], $R$ and $\nabla$ are associated with a metric tensor $M$, and the tensor $\tau_{i}^{jk}\equiv\nabla_{\bullet}^{*}N^{*}$ is given by a covariant derivative of a dual induced tensor,
@@
The proof of proposition 1 follows straightforwardly from [11, 12] keeping track of the extra terms. For the purpose of this paper, we skip the coordinate-invariant version of the equations. The coordinate-invariant approach will be discussed elsewhere.